the sum of the page numbers on facing pages of a book is 157, what are the page numbers?
Answer by checkley77(12844)

x+(x+1)=157
2x+1=157
2x=157-1
2x=156
x=156/2
x=78 for the left hand page number.
78+1=79 for the right hand page number.
Proof:
78+79=157
157=157